\begin{thebibliography}{72}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi
\expandafter\ifx\csname bibnamefont\endcsname\relax
  \def\bibnamefont#1{#1}\fi
\expandafter\ifx\csname bibfnamefont\endcsname\relax
  \def\bibfnamefont#1{#1}\fi
\expandafter\ifx\csname citenamefont\endcsname\relax
  \def\citenamefont#1{#1}\fi
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[{\citenamefont{Oldham and Spanier}(2006)}]{Oldham2006}
\bibinfo{author}{\bibfnamefont{K.~B.} \bibnamefont{Oldham}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Spanier}},
  \emph{\bibinfo{title}{The Fractional Calculus: Theory and Applications of
  Differentiation and Integration to Arbitrary Order}}, Dover Books on
  Mathematics (\bibinfo{publisher}{Dover Publications}, \bibinfo{year}{2006}).

\bibitem[{\citenamefont{Uchaikin}(2008)}]{Uchaikin2008500}
\bibinfo{author}{\bibfnamefont{V.}~\bibnamefont{Uchaikin}},
  \emph{\bibinfo{title}{Fractional Derivatives Method}}
  (\bibinfo{publisher}{Artishock-Press}, \bibinfo{address}{Uljanovsk},
  \bibinfo{year}{2008}).

\bibitem[{\citenamefont{Sahimi}(1998)}]{Sahimi1998213}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Sahimi}},
  \bibinfo{journal}{Physics Reports} \textbf{\bibinfo{volume}{306}},
  \bibinfo{pages}{213} (\bibinfo{year}{1998}).

\bibitem[{\citenamefont{Koro\u{s}ak et~al.}(2007)\citenamefont{Koro\u{s}ak,
  Cvikl, Kramer, Jecl, and Prapotnik}}]{Korosak20071}
\bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Koro\u{s}ak}},
  \bibinfo{author}{\bibfnamefont{B.}~\bibnamefont{Cvikl}},
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Kramer}},
  \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Jecl}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Prapotnik}},
  \bibinfo{journal}{Journal of Contaminant Hydrology}
  \textbf{\bibinfo{volume}{92}}, \bibinfo{pages}{1} (\bibinfo{year}{2007}).

\bibitem[{\citenamefont{Hobbie and Roth}(2007)}]{Hobbie2007}
\bibinfo{author}{\bibfnamefont{R.~K.} \bibnamefont{Hobbie}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{B.~J.} \bibnamefont{Roth}},
  \emph{\bibinfo{title}{Intermediate Physics for Medicine and Biology}}
  (\bibinfo{publisher}{Springer}, \bibinfo{address}{New York},
  \bibinfo{year}{2007}).

\bibitem[{\citenamefont{Metzler and Klafter}(2000)}]{Metzler20001}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Metzler}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Klafter}},
  \bibinfo{journal}{Physics Reports} \textbf{\bibinfo{volume}{339}},
  \bibinfo{pages}{1} (\bibinfo{year}{2000}).

\bibitem[{\citenamefont{Hilfer}(1995)}]{Hilfer19951475}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Hilfer}},
  \bibinfo{journal}{Chaos, Solitons \& Fractals} \textbf{\bibinfo{volume}{5}},
  \bibinfo{pages}{1475 } (\bibinfo{year}{1995}).

\bibitem[{\citenamefont{Hilfer}(2000{\natexlab{a}})}]{Hilfer20003914}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Hilfer}}, \bibinfo{journal}{The
  Journal of Physical Chemistry B} \textbf{\bibinfo{volume}{104}},
  \bibinfo{pages}{3914} (\bibinfo{year}{2000}{\natexlab{a}}).

\bibitem[{\citenamefont{Hilfer}(2000{\natexlab{b}})}]{Hilfer2000}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Hilfer}},
  \emph{\bibinfo{title}{Fractional Time Evolution}} (\bibinfo{publisher}{World
  Scientific}, \bibinfo{address}{Singapore, New Jersey, London, Hong Kong},
  \bibinfo{year}{2000}{\natexlab{b}}), chap.~\bibinfo{chapter}{II}, pp.
  \bibinfo{pages}{87--130}.

\bibitem[{\citenamefont{Koszto\l{}owicz
  et~al.}(2005)\citenamefont{Koszto\l{}owicz, Dworecki, and
  Mr\'owczy\ifmmode~\acute{n}\else \'{n}\fi{}ski}}]{Kosztolowicz2005041105}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Koszto\l{}owicz}},
  \bibinfo{author}{\bibfnamefont{K.}~\bibnamefont{Dworecki}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{S.}~\bibnamefont{Mr\'owczy\ifmmode~\acute{n}\else
  \'{n}\fi{}ski}}, \bibinfo{journal}{Phys. Rev. E}
  \textbf{\bibinfo{volume}{71}}, \bibinfo{pages}{041105}
  (\bibinfo{year}{2005}).

\bibitem[{\citenamefont{Koszto\l{}owicz}(2015)}]{Kosztolowicz2015P10021}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Koszto\l{}owicz}},
  \bibinfo{journal}{Journal of Statistical Mechanics: Theory and Experiment}
  \textbf{\bibinfo{volume}{2015}}, \bibinfo{pages}{P10021}
  (\bibinfo{year}{2015}).

\bibitem[{\citenamefont{Bisquert et~al.}(2000)\citenamefont{Bisquert,
  Garcia-Belmonte, Fabregat-Santiago, Ferriols, Bogdanoff, and
  Pereira}}]{Bisquert20002287}
\bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Bisquert}},
  \bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{Garcia-Belmonte}},
  \bibinfo{author}{\bibfnamefont{F.}~\bibnamefont{Fabregat-Santiago}},
  \bibinfo{author}{\bibfnamefont{N.~S.} \bibnamefont{Ferriols}},
  \bibinfo{author}{\bibfnamefont{P.}~\bibnamefont{Bogdanoff}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{E.~C.}
  \bibnamefont{Pereira}}, \bibinfo{journal}{The Journal of Physical Chemistry
  B} \textbf{\bibinfo{volume}{104}}, \bibinfo{pages}{2287}
  (\bibinfo{year}{2000}).

\bibitem[{\citenamefont{Bisquert and Compte}(2001)}]{Bisquert2001112}
\bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Bisquert}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Compte}},
  \bibinfo{journal}{Journal of Electroanalytical Chemistry}
  \textbf{\bibinfo{volume}{499}}, \bibinfo{pages}{112} (\bibinfo{year}{2001}).

\bibitem[{\citenamefont{Koszto\l{}owicz and
  Lewandowska}(2009)}]{Kosztolowicz2009055004}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Koszto\l{}owicz}}
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{K.~D.}
  \bibnamefont{Lewandowska}}, \bibinfo{journal}{Journal of Physics A:
  Mathematical and Theoretical} \textbf{\bibinfo{volume}{42}},
  \bibinfo{pages}{055004} (\bibinfo{year}{2009}).

\bibitem[{\citenamefont{Pyanylo et~al.}(2014)\citenamefont{Pyanylo, Prytula,
  Prytula, and Lopuh}}]{Pyanylo201484}
\bibinfo{author}{\bibfnamefont{Y.}~\bibnamefont{Pyanylo}},
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Prytula}},
  \bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Prytula}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Lopuh}},
  \bibinfo{journal}{Mathematical Modeling and Computing}
  \textbf{\bibinfo{volume}{1}}, \bibinfo{pages}{84} (\bibinfo{year}{2014}).

\bibitem[{\citenamefont{Berkowitz and Scher}(1998)}]{Berkowitz19985858}
\bibinfo{author}{\bibfnamefont{B.}~\bibnamefont{Berkowitz}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{H.}~\bibnamefont{Scher}},
  \bibinfo{journal}{Phys. Rev. E} \textbf{\bibinfo{volume}{57}},
  \bibinfo{pages}{5858} (\bibinfo{year}{1998}).

\bibitem[{\citenamefont{Bouchaud and Georges}(1990)}]{Bouchaud1990127}
\bibinfo{author}{\bibfnamefont{J.-P.} \bibnamefont{Bouchaud}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Georges}},
  \bibinfo{journal}{Physics Reports} \textbf{\bibinfo{volume}{195}},
  \bibinfo{pages}{127} (\bibinfo{year}{1990}).

\bibitem[{\citenamefont{Balescu}(1995)}]{Balescu19954807}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Balescu}},
  \bibinfo{journal}{Phys. Rev. E} \textbf{\bibinfo{volume}{51}},
  \bibinfo{pages}{4807} (\bibinfo{year}{1995}).

\bibitem[{\citenamefont{Tribeche and Shukla}(2011)}]{Tribeche2011103702}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Tribeche}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{P.~K.} \bibnamefont{Shukla}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{18}},
  \bibinfo{pages}{103702} (\bibinfo{year}{2011}).

\bibitem[{\citenamefont{Gong and Du}(2012)}]{Gong2012023704}
\bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Gong}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Du}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{19}},
  \bibinfo{pages}{023704} (\bibinfo{year}{2012}).

\bibitem[{\citenamefont{Carreras et~al.}(2001)\citenamefont{Carreras, Lynch,
  and Zaslavsky}}]{Carreras20015096}
\bibinfo{author}{\bibfnamefont{B.~A.} \bibnamefont{Carreras}},
  \bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Lynch}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{G.~M.} \bibnamefont{Zaslavsky}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{8}},
  \bibinfo{pages}{5096} (\bibinfo{year}{2001}).

\bibitem[{\citenamefont{Tarasov}(2005{\natexlab{a}})}]{Tarasov2005082106}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{12}},
  \bibinfo{pages}{082106} (\bibinfo{year}{2005}{\natexlab{a}}).

\bibitem[{\citenamefont{Tarasov}(2006{\natexlab{a}})}]{Tarasov2006052107}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{13}},
  \bibinfo{pages}{052107} (\bibinfo{year}{2006}{\natexlab{a}}).

\bibitem[{\citenamefont{Monin}(1955)}]{Monin1955256}
\bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Monin}}, \bibinfo{journal}{DAN
  SSSR, ser. geofiz.} \textbf{\bibinfo{volume}{2}}, \bibinfo{pages}{256}
  (\bibinfo{year}{1955}).

\bibitem[{\citenamefont{Klimontovich}(2002)}]{Klimontovich2002}
\bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Klimontovich}},
  \emph{\bibinfo{title}{Vvedenie v fiziku otkrytyh sistem}}
  (\bibinfo{publisher}{Moskva Janus}, \bibinfo{year}{2002}).

\bibitem[{\citenamefont{Zaslavsky}(2002)}]{Zaslavsky2002461}
\bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{Zaslavsky}},
  \bibinfo{journal}{Physics Reports} \textbf{\bibinfo{volume}{371}},
  \bibinfo{pages}{461} (\bibinfo{year}{2002}).

\bibitem[{\citenamefont{Zaslavsky}(1994)}]{Zaslavsky1994110}
\bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{Zaslavsky}},
  \bibinfo{journal}{Physica D: Nonlinear Phenomena}
  \textbf{\bibinfo{volume}{76}}, \bibinfo{pages}{110} (\bibinfo{year}{1994}).

\bibitem[{\citenamefont{Saichev and Zaslavsky}(1997)}]{Saichev1997753}
\bibinfo{author}{\bibfnamefont{A.~I.} \bibnamefont{Saichev}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{G.~M.} \bibnamefont{Zaslavsky}},
  \bibinfo{journal}{Chaos} \textbf{\bibinfo{volume}{7}}, \bibinfo{pages}{753}
  (\bibinfo{year}{1997}).

\bibitem[{\citenamefont{Zaslavsky and Edelman}(2004)}]{Zaslavsky2004128}
\bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{Zaslavsky}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Edelman}},
  \bibinfo{journal}{Physica D: Nonlinear Phenomena}
  \textbf{\bibinfo{volume}{193}}, \bibinfo{pages}{128} (\bibinfo{year}{2004}).

\bibitem[{\citenamefont{Nigmatullin}(2006)}]{Nigmatullin2006282}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Nigmatullin}},
  \bibinfo{journal}{Physica A: Statistical Mechanics and its Applications}
  \textbf{\bibinfo{volume}{363}}, \bibinfo{pages}{282} (\bibinfo{year}{2006}).

\bibitem[{\citenamefont{Chechkin et~al.}(2002)\citenamefont{Chechkin, Gonchar,
  and Szyd{\l}owski}}]{Chechkin200278}
\bibinfo{author}{\bibfnamefont{A.~V.} \bibnamefont{Chechkin}},
  \bibinfo{author}{\bibfnamefont{V.~Y.} \bibnamefont{Gonchar}},
  \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Szyd{\l}owski}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{9}},
  \bibinfo{pages}{78} (\bibinfo{year}{2002}).

\bibitem[{\citenamefont{Gafiychuk and Datsko}(2007)}]{Gafiychuk2007055201}
\bibinfo{author}{\bibfnamefont{V.~V.} \bibnamefont{Gafiychuk}}
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{B.~Y.}
  \bibnamefont{Datsko}}, \bibinfo{journal}{Phys. Rev. E}
  \textbf{\bibinfo{volume}{75}}, \bibinfo{pages}{055201}
  (\bibinfo{year}{2007}).

\bibitem[{\citenamefont{Koszto\l{}owicz and
  Lewandowska}(2008)}]{Kosztolowicz2008066103}
\bibinfo{author}{\bibfnamefont{T.}~\bibnamefont{Koszto\l{}owicz}}
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{K.~D.}
  \bibnamefont{Lewandowska}}, \bibinfo{journal}{Phys. Rev. E}
  \textbf{\bibinfo{volume}{78}}, \bibinfo{pages}{066103}
  (\bibinfo{year}{2008}).

\bibitem[{\citenamefont{Shkilev}(2013)}]{Shkilev20131066}
\bibinfo{author}{\bibfnamefont{V.~P.} \bibnamefont{Shkilev}},
  \bibinfo{journal}{Journal of Experimental and Theoretical Physics}
  \textbf{\bibinfo{volume}{117}}, \bibinfo{pages}{1066} (\bibinfo{year}{2013}).

\bibitem[{\citenamefont{Laskin}(2000{\natexlab{a}})}]{Laskin2000780}
\bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Laskin}},
  \bibinfo{journal}{Chaos} \textbf{\bibinfo{volume}{10}}, \bibinfo{pages}{780}
  (\bibinfo{year}{2000}{\natexlab{a}}).

\bibitem[{\citenamefont{Laskin}(2000{\natexlab{b}})}]{Laskin20003135}
\bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Laskin}},
  \bibinfo{journal}{Phys. Rev. E} \textbf{\bibinfo{volume}{62}},
  \bibinfo{pages}{3135} (\bibinfo{year}{2000}{\natexlab{b}}).

\bibitem[{\citenamefont{Laskin}(2000{\natexlab{c}})}]{Laskin2000298}
\bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Laskin}},
  \bibinfo{journal}{Physics Letters A} \textbf{\bibinfo{volume}{268}},
  \bibinfo{pages}{298 } (\bibinfo{year}{2000}{\natexlab{c}}).

\bibitem[{\citenamefont{Laskin}(2002)}]{Laskin2002056108}
\bibinfo{author}{\bibfnamefont{N.}~\bibnamefont{Laskin}},
  \bibinfo{journal}{Phys. Rev. E} \textbf{\bibinfo{volume}{66}},
  \bibinfo{pages}{056108} (\bibinfo{year}{2002}).

\bibitem[{\citenamefont{Naber}(2004)}]{Naber20043339}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Naber}},
  \bibinfo{journal}{Journal of Mathematical Physics}
  \textbf{\bibinfo{volume}{45}}, \bibinfo{pages}{3339} (\bibinfo{year}{2004}).

\bibitem[{\citenamefont{Uchaikin}(2013)}]{Uchaikin20131074}
\bibinfo{author}{\bibfnamefont{V.~V.} \bibnamefont{Uchaikin}},
  \bibinfo{journal}{Physics-Uspekhi} \textbf{\bibinfo{volume}{56}},
  \bibinfo{pages}{1074} (\bibinfo{year}{2013}).

\bibitem[{\citenamefont{O'Shaughnessy and
  Procaccia}(1985)}]{O'Shaughnessy1985455}
\bibinfo{author}{\bibfnamefont{B.}~\bibnamefont{O'Shaughnessy}}
  \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{I.}~\bibnamefont{Procaccia}},
  \bibinfo{journal}{Phys. Rev. Lett.} \textbf{\bibinfo{volume}{54}},
  \bibinfo{pages}{455} (\bibinfo{year}{1985}).

\bibitem[{\citenamefont{Mandelbrot}(1982)}]{Mandelbrot1982}
\bibinfo{author}{\bibfnamefont{B.~B.} \bibnamefont{Mandelbrot}},
  \emph{\bibinfo{title}{The fractal geometry of nature}}
  (\bibinfo{publisher}{W. H. Freeman and Company}, \bibinfo{year}{1982}),
  \bibinfo{edition}{1st} ed.

\bibitem[{\citenamefont{Metzler et~al.}(1999)\citenamefont{Metzler, Barkai, and
  Klafter}}]{Metzler1999431}
\bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Metzler}},
  \bibinfo{author}{\bibfnamefont{E.}~\bibnamefont{Barkai}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Klafter}},
  \bibinfo{journal}{Europhysics Letters} \textbf{\bibinfo{volume}{46}},
  \bibinfo{pages}{431} (\bibinfo{year}{1999}).

\bibitem[{\citenamefont{Essex et~al.}(2000)\citenamefont{Essex, Schulzky,
  Franz, and Hoffmann}}]{Essex2000299}
\bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Essex}},
  \bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Schulzky}},
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Franz}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{K.~H.} \bibnamefont{Hoffmann}},
  \bibinfo{journal}{Physica A: Statistical Mechanics and its Applications}
  \textbf{\bibinfo{volume}{284}}, \bibinfo{pages}{299} (\bibinfo{year}{2000}).

\bibitem[{\citenamefont{Tsallis}(2001)}]{Tsallis2001}
\bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Tsallis}},
  \emph{\bibinfo{title}{Nonextensive Statistical Mechanics and Its
  Applications}} (\bibinfo{publisher}{Springer-Verlag Berlin Heidelberg},
  \bibinfo{year}{2001}), Lecture Notes in Physics 560, \bibinfo{edition}{1st}
  ed.

\bibitem[{\citenamefont{Gell-Mann and Tsallis}(2004)}]{Gell-Mann2004}
\bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Gell-Mann}} \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{C.}~\bibnamefont{Tsallis}},
  \emph{\bibinfo{title}{Nonextensive entropy: Interdisciplinary applications}},
  Santa Fe Institute Studies on the Sciences of Complexity
  (\bibinfo{publisher}{Oxford University Press, USA}, \bibinfo{year}{2004}).

\bibitem[{\citenamefont{Vasconcellos et~al.}(2006)\citenamefont{Vasconcellos,
  Galv\~{a}o Ramos, Gorenstein, Kleinke, Souza~Cruz, and
  Luzzi}}]{Vasconcellos20064821}
\bibinfo{author}{\bibfnamefont{A.~R.} \bibnamefont{Vasconcellos}},
  \bibinfo{author}{\bibfnamefont{J.}~\bibnamefont{Galv\~{a}o Ramos}},
  \bibinfo{author}{\bibfnamefont{A.}~\bibnamefont{Gorenstein}},
  \bibinfo{author}{\bibfnamefont{M.~U.} \bibnamefont{Kleinke}},
  \bibinfo{author}{\bibfnamefont{T.~G.} \bibnamefont{Souza~Cruz}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Luzzi}},
  \bibinfo{journal}{International Journal of Modern Physics B}
  \textbf{\bibinfo{volume}{20}}, \bibinfo{pages}{4821} (\bibinfo{year}{2006}).

\bibitem[{\citenamefont{Uchaikin}(2003)}]{Uchaikin2003810}
\bibinfo{author}{\bibfnamefont{V.~V.} \bibnamefont{Uchaikin}},
  \bibinfo{journal}{Journal of Experimental and Theoretical Physics}
  \textbf{\bibinfo{volume}{97}}, \bibinfo{pages}{810} (\bibinfo{year}{2003}).

\bibitem[{\citenamefont{Stanislavsky}(2004)}]{Stanislavsky2004418}
\bibinfo{author}{\bibfnamefont{A.~A.} \bibnamefont{Stanislavsky}},
  \bibinfo{journal}{Theoretical and Mathematical Physics}
  \textbf{\bibinfo{volume}{138}}, \bibinfo{pages}{418} (\bibinfo{year}{2004}).

\bibitem[{\citenamefont{Tarasov}(2004)}]{Tarasov2004123}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Chaos} \textbf{\bibinfo{volume}{14}}, \bibinfo{pages}{123}
  (\bibinfo{year}{2004}).

\bibitem[{\citenamefont{Tarasov}(2005{\natexlab{b}})}]{Tarasov200517}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Journal of Physics: Conference Series}
  \textbf{\bibinfo{volume}{7}}, \bibinfo{pages}{17}
  (\bibinfo{year}{2005}{\natexlab{b}}).

\bibitem[{\citenamefont{Tarasov}(2005{\natexlab{c}})}]{Tarasov2005011102}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Phys. Rev. E} \textbf{\bibinfo{volume}{71}},
  \bibinfo{pages}{011102} (\bibinfo{year}{2005}{\natexlab{c}}).

\bibitem[{\citenamefont{Tarasov}(2006{\natexlab{b}})}]{Tarasov2006033108}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Chaos} \textbf{\bibinfo{volume}{16}},
  \bibinfo{pages}{033108} (\bibinfo{year}{2006}{\natexlab{b}}).

\bibitem[{\citenamefont{Tarasov}(2006{\natexlab{c}})}]{Tarasov2006341}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{International Journal of Modern Physics B}
  \textbf{\bibinfo{volume}{20}}, \bibinfo{pages}{341}
  (\bibinfo{year}{2006}{\natexlab{c}}).

\bibitem[{\citenamefont{Tarasov}(2013{\natexlab{a}})}]{Tarasov2013663}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Nonlinear Dynamics} \textbf{\bibinfo{volume}{71}},
  \bibinfo{pages}{663} (\bibinfo{year}{2013}{\natexlab{a}}).

\bibitem[{\citenamefont{Tarasov}(2008)}]{Tarasov20082984}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Physics Letters A} \textbf{\bibinfo{volume}{372}},
  \bibinfo{pages}{2984 } (\bibinfo{year}{2008}).

\bibitem[{\citenamefont{Tarasov}(2005{\natexlab{d}})}]{Tarasov2005286}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Annals of Physics} \textbf{\bibinfo{volume}{318}},
  \bibinfo{pages}{286 } (\bibinfo{year}{2005}{\natexlab{d}}).

\bibitem[{\citenamefont{Tarasov}(2007{\natexlab{a}})}]{Tarasov2007237}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Modern Physics Letters B} \textbf{\bibinfo{volume}{21}},
  \bibinfo{pages}{237} (\bibinfo{year}{2007}{\natexlab{a}}).

\bibitem[{\citenamefont{Tarasov}(2007{\natexlab{b}})}]{Tarasov2007163}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Modern Physics Letters B} \textbf{\bibinfo{volume}{21}},
  \bibinfo{pages}{163} (\bibinfo{year}{2007}{\natexlab{b}}).

\bibitem[{\citenamefont{Tarasov}(2009)}]{Tarasov2009179}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Theoretical and Mathematical Physics}
  \textbf{\bibinfo{volume}{158}}, \bibinfo{pages}{179} (\bibinfo{year}{2009}).

\bibitem[{\citenamefont{Tarasov}(2012)}]{Tarasov20121719}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Annals of Physics} \textbf{\bibinfo{volume}{327}},
  \bibinfo{pages}{1719} (\bibinfo{year}{2012}).

\bibitem[{\citenamefont{Tarasov}(2013{\natexlab{b}})}]{Tarasov2013102110}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \bibinfo{journal}{Physics of Plasmas} \textbf{\bibinfo{volume}{20}},
  \bibinfo{pages}{102110} (\bibinfo{year}{2013}{\natexlab{b}}).

\bibitem[{\citenamefont{Tarasov}(2010)}]{Tarasov2010}
\bibinfo{author}{\bibfnamefont{V.~E.} \bibnamefont{Tarasov}},
  \emph{\bibinfo{title}{Fractional Dynamics: Applications of Fractional
  Calculus to Dynamics of Particles, Fields and Media}}, Nonlinear Physical
  Science (\bibinfo{publisher}{Springer Berlin Heidelberg},
  \bibinfo{year}{2010}), \bibinfo{edition}{1st} ed.

\bibitem[{\citenamefont{Kobelev}(2000)}]{Kobelev20000002002}
\bibinfo{author}{\bibfnamefont{L.~Y.} \bibnamefont{Kobelev}},
  \emph{\bibinfo{title}{The multifractal time and irreversibility in dynamic
  systems}} (\bibinfo{year}{2000}), \eprint{arXiv:physics/0002002}.

\bibitem[{\citenamefont{Kobelev et~al.}(2000)\citenamefont{Kobelev, Kobelev,
  and Romanov}}]{Kobelev2000194}
\bibinfo{author}{\bibfnamefont{Y.~L.} \bibnamefont{Kobelev}},
  \bibinfo{author}{\bibfnamefont{L.~Y.} \bibnamefont{Kobelev}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{E.~P.}
  \bibnamefont{Romanov}}, \bibinfo{journal}{Doklady Physics}
  \textbf{\bibinfo{volume}{45}}, \bibinfo{pages}{194} (\bibinfo{year}{2000}).

\bibitem[{\citenamefont{Kobelev et~al.}(2002)\citenamefont{Kobelev, Kobelev,
  Kobelev, and Romanov}}]{Kobelev2002580}
\bibinfo{author}{\bibfnamefont{Y.~L.} \bibnamefont{Kobelev}},
  \bibinfo{author}{\bibfnamefont{L.~Y.} \bibnamefont{Kobelev}},
  \bibinfo{author}{\bibfnamefont{V.~L.} \bibnamefont{Kobelev}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{E.~P.}
  \bibnamefont{Romanov}}, \bibinfo{journal}{Doklady Physics}
  \textbf{\bibinfo{volume}{47}}, \bibinfo{pages}{580} (\bibinfo{year}{2002}).

\bibitem[{\citenamefont{Lukashchuk}(2013)}]{Lukashchuk2013740}
\bibinfo{author}{\bibfnamefont{S.~Y.} \bibnamefont{Lukashchuk}},
  \bibinfo{journal}{Central European Journal of Physics}
  \textbf{\bibinfo{volume}{11}}, \bibinfo{pages}{740} (\bibinfo{year}{2013}).

\bibitem[{\citenamefont{Zubarev}(1981)}]{Zubarev19811509}
\bibinfo{author}{\bibfnamefont{D.~N.} \bibnamefont{Zubarev}},
  \bibinfo{journal}{Journal of Soviet Mathematics}
  \textbf{\bibinfo{volume}{16}}, \bibinfo{pages}{1509} (\bibinfo{year}{1981}).

\bibitem[{\citenamefont{Zubarev
  et~al.}(2002{\natexlab{a}})\citenamefont{Zubarev, Morozov, and
  R\"{o}pke}}]{Zubarev20021}
\bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Zubarev}},
  \bibinfo{author}{\bibfnamefont{V.}~\bibnamefont{Morozov}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{R\"{o}pke}},
  \emph{\bibinfo{title}{Statistical mechanics of nonequilibrium processes}},
  vol.~\bibinfo{volume}{1} (\bibinfo{publisher}{Fizmatlit},
  \bibinfo{year}{2002}{\natexlab{a}}).

\bibitem[{\citenamefont{Zubarev
  et~al.}(2002{\natexlab{b}})\citenamefont{Zubarev, Morozov, and
  R\"{o}pke}}]{Zubarev20022}
\bibinfo{author}{\bibfnamefont{D.}~\bibnamefont{Zubarev}},
  \bibinfo{author}{\bibfnamefont{V.}~\bibnamefont{Morozov}}, \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{G.}~\bibnamefont{R\"{o}pke}},
  \emph{\bibinfo{title}{Statistical mechanics of nonequilibrium processes}},
  vol.~\bibinfo{volume}{2} (\bibinfo{publisher}{Fizmatlit},
  \bibinfo{year}{2002}{\natexlab{b}}).

\bibitem[{\citenamefont{Markiv et~al.}(2011)\citenamefont{Markiv, Tokarchuk,
  Kostrobij, and Tokarchuk}}]{Markiv2011785}
\bibinfo{author}{\bibfnamefont{B.}~\bibnamefont{Markiv}},
  \bibinfo{author}{\bibfnamefont{R.}~\bibnamefont{Tokarchuk}},
  \bibinfo{author}{\bibfnamefont{P.}~\bibnamefont{Kostrobij}},
  \bibnamefont{and}
  \bibinfo{author}{\bibfnamefont{M.}~\bibnamefont{Tokarchuk}},
  \bibinfo{journal}{Physica A: Statistical Mechanics and its Applications}
  \textbf{\bibinfo{volume}{390}}, \bibinfo{pages}{785} (\bibinfo{year}{2011}).

\bibitem[{\citenamefont{Samko et~al.}(1993)\citenamefont{Samko, Kilbas, and
  Marichev}}]{Samko1993}
\bibinfo{author}{\bibfnamefont{S.~G.} \bibnamefont{Samko}},
  \bibinfo{author}{\bibfnamefont{A.~A.} \bibnamefont{Kilbas}},
  \bibnamefont{and} \bibinfo{author}{\bibfnamefont{O.~I.}
  \bibnamefont{Marichev}}, \emph{\bibinfo{title}{Fractional Integrals and
  Derivatives: Theory and Applications}} (\bibinfo{publisher}{Gordon and Breach
  Science Publishers}, \bibinfo{year}{1993}), \bibinfo{edition}{1st} ed.

\end{thebibliography}

\documentclass[aps,pre,showpacs,showkeys,groupedaddress,preprint]{revtex4-1}  
\usepackage[cp1251]{inputenc}
\usepackage{graphicx}
\usepackage[english]{babel}
\usepackage{amsmath}
\begin{document}

\title{Generalized diffusion equation with fractional derivatives within Renyi statistics}

\author{P. Kostrobij$^{1}$, B. Markovych$^{1}$, O. Viznovych$^{1}$, M. Tokarchuk$^{1,2}$}

\affiliation{$^1$Lviv Polytechnic National University, 12 Bandera Str., 79013 Lviv, Ukraine\\
$^2$Institute for Condensed Matter Physics of NAS of Ukraine, 1 Svientsitskii Str.,
79011 Lviv, Ukraine}


\begin{abstract}
 By using the Zubarev nonequilibrium statistical operator method,
 and the Liouville equation with fractional derivatives,
 a generalized diffusion equation with fractional derivatives is obtained within the Renyi statistics.
 Averaging in generalized diffusion coefficient is performed with a power distribution with the Renyi parameter~$q$.
\end{abstract}

\keywords{generalized diffusion equation, nonequilibrium statistical operator, Renyi statistics, anomalous diffusion}

\maketitle

 \section{Introduction}

 The fractional derivatives and integrals~\cite{Oldham2006} are widely used
 to study anomalous diffusion in porous media~\cite{Uchaikin2008500,Sahimi1998213,Korosak20071,Hobbie2007,Metzler20001,Hilfer19951475,Hilfer20003914,Hilfer2000,Kosztolowicz2005041105,Kosztolowicz2015P10021,Bisquert20002287,Bisquert2001112,Kosztolowicz2009055004,Pyanylo201484},
 in disordered systems~\cite{Berkowitz19985858,Bouchaud1990127},
 in plasma physics~\cite{Balescu19954807,Tribeche2011103702,Gong2012023704,Carreras20015096,Tarasov2005082106,Tarasov2006052107},
 in turbulent~\cite{Monin1955256,Klimontovich2002,Zaslavsky2002461},
 kinetic and reaction-diffusion processes~\cite{Zaslavsky2002461,Zaslavsky1994110,Saichev1997753,Zaslavsky2004128,Nigmatullin2006282,Chechkin200278,Gafiychuk2007055201,Kosztolowicz2008066103,Shkilev20131066},
 in quantum mechanics~\cite{Laskin2000780,Laskin20003135,Laskin2000298,Laskin2002056108,Naber20043339}, etc.~\cite{Uchaikin2008500,Uchaikin20131074}

 Experimental data on the different processes of anomalous diffusion show that
 not only the law distribution,
 but also form of diffusion package is significantly different from the normal diffusion~\cite{Uchaikin2008500,Bouchaud1990127,Zaslavsky2002461,Uchaikin20131074}.
 Approaches with variable diffusion coefficients \cite{O'Shaughnessy1985455}, on the basis of a degree correlation of fractional order \cite{Mandelbrot1982}, fractional derivatives \cite{Monin1955256,Klimontovich2002,Zaslavsky2002461}, generalized Fokker-Planck equation \cite{Metzler20001,Zaslavsky2002461,Metzler1999431}, generalization of statistical mechanics (extensive and non-extensive) based on the Tsallis \cite{Essex2000299,Tsallis2001,Gell-Mann2004} and Renyj~\cite{Essex2000299,Vasconcellos20064821} entropy, and others were developed to describe the anomalous diffusion in different physical and chemical systems.
 Conducted researches show that
 mathematical basis of anomalous diffusion is equation with fractional derivatives \cite{Uchaikin2008500,Zaslavsky2002461}.
 In particular,
 during the study of three-dimensional models of anomalous diffusion~\cite{Uchaikin2008500,Uchaikin20131074,Uchaikin2003810},
 the basic equation of anomalous diffusion is derived
 from the general principles of the stochastic theory of random processes
 based on the Chapman-Kolmogorov integral equations for transition probability.
 Solutions of these equations form a new class of distributions, called fractional stable distributions.
 These distributions are solutions of partial differential equations of fractional order.
 These equations are generalization of usual diffusion equation to the case of anomalous diffusion.
 A partial case of the fractional stable distribution is the Gaussian distribution,
 which corresponds to the normal diffusion.
 It is important to note that the equations for anomalous diffusion with fractional derivatives
 contain diffusion coefficient, which is  a constant in time and space.
 On the other hand, the diffusion coefficients are related to time correlation functions
 (the Green-Kubo relations)
 containing diffusion transfer mechanisms from the perspective of nonequilibrium statistical mechanics.

 Currently,
 together with the phenomenological approach for constructing of the Fokker–Planck equation,
 the diffusion equation and its generalization --- the Cattaneo equation with fractional derivatives,
 there are two methods of constructing such equations,
 namely,
 (1)~probabilistic method,
 based on the Chapman-Kolmogorov equation in stochastic theory of random processes~\cite{Uchaikin2008500,Zaslavsky2002461,Stanislavsky2004418},
 and (2)~statistical method,
 based on the Liouville equation with fractional
 derivatives~\cite{Tarasov2004123,Tarasov200517,Tarasov2005011102,Tarasov2006033108,Tarasov2006341,Tarasov2013663,Tarasov20082984,Tarasov2005286,Tarasov2007237,Tarasov2007163,Tarasov2009179,Tarasov20121719,Tarasov2013102110,Tarasov2010}.
 In particular,
 by using this method,
 the BBGKY hierarchy equations with fractional derivatives~\cite{Tarasov200517,Tarasov2005011102,Tarasov2007237},
 transport equation,
 diffusion equation,
 and the Heisenberg equation with fractional derivatives~\cite{Tarasov2006341,Tarasov2013663,Tarasov20082984} are obtained.
 This approach is formulated for non-Hamiltonian systems.
 If the Helmholtz conditions for the coordinate and momentum derivatives are fulfilled,
 the Hamiltonian systems with the time-reversible Liouville equation with fractional derivatives are obtained
 from non-Hamiltonian systems.
 In Ref.~\cite{Kobelev20000002002},
 time-irreversible equations of motion of Hamilton and Liouville
 for dynamic of classical particles in space with multifractal time are offered.
 By using the definition of fractional derivative and the Riemann-Liouville integral,
 the time-irreversible Liouville equation with fractional derivatives
 (where the time is given on multifractal sets with fractional dimensions) is obtained.
 In Refs.~\cite{Kobelev2000194,Kobelev2002580},
 kinetic equation for systems with fractal structure
 (in particular, for description of diffusion processes in space of coordinates and momenta) is obtained
 within the Klimontovich approach.
 A similar approach for constructing of time fractional generalization
 for the Liouville equation and the Zwanzig equation (within projection formalism) is proposed in Ref.~\cite{Lukashchuk2013740}.

 In the present work,
 by using the Zubarev nonequilibrium statistical operator method~\cite{Zubarev19811509,Zubarev20021,Zubarev20022,Markiv2011785} and the maximum entropy principle for the Renyi entropies,
 we consider a way of obtaining generalized (non-Markovian) diffusion equation with fractional derivatives.
 Using the Liouville equation with fractional derivatives proposed by Tarasov In Refs.~\cite{Tarasov2004123,Tarasov200517,Tarasov2005011102,Tarasov2006033108}
 is an important and fundamental step for obtaining this equation.
 In the second section,
 by using the Zubarev nonequilibrium statistical operator method and
 the maximum entropy principle for the Renyi entropies,
 we found a solution of the Liouville equation with fractional derivatives at a selected set of observed variables.
 In the third section, we chose nonequilibrium average values of particle density as a parameter of reduced description,
 and then we received a generalized (non-Markovian) diffusion equation with fractional derivatives.
 
 \newpage

\section{Liouville equation with fractional derivatives for classical system of particles}

 We use the Liouville equation with fractional derivatives
 obtained by Tarasov in Refs.~\cite{Tarasov2004123,Tarasov200517,Tarasov2005011102,Tarasov2006033108}
 for non-equilibrium particle function $\rho(x^{N};t)$ of classical system:
 \begin{equation}\label{eq:2.1}
  \frac{\partial }{\partial t}\rho(x^{N};t)+\sum^{N}_{j=1}D^{\alpha}_{\vec{r}_{j}}
  \left(\rho(x^{N};t)\vec{v}_{j}\right)+\sum^{N}_{j=1}D^{\alpha}_{\vec{p}_{j}}
  \left(\rho(x^{N};t)\vec{F}_{j}\right)=0,
 \end{equation}
 where $x^{N}=x_{1},\ldots,x_{N}$,
 $x_{j}=\{\vec{r}_{j}, \vec{p}_{j}\}$ are the phase variables (coordinate and momentum) of $j$-th particle,
 $\vec{v}_{j}$ is the fields of velocity,
 $\vec{F}_{j}$ is the force field acting on $j$-th particle.
 \begin{equation}\label{eq:2.2}
  D^{\alpha}_{x} f(x)=\frac{1}{\Gamma (n-\alpha)} \int^{x}_{0} \frac{f^{n}(z)}{(x-z)^{\alpha +1-n}} dz
 \end{equation}
 is the Riemann-Liouville fractional derivative~\cite{Oldham2006,Samko1993},
 $n-1<\alpha<n$,
 $f^{n}(z)=\frac{d^{n}}{dz^{n}}f(z)$.
 In the general case we have
 \[
  D^{\alpha}_{\vec{r}_{j}}\left(\rho(x^{N};t)\vec{F}_{j}\right)
  \neq
  \rho(x^{N};t)D^{\alpha}_{\vec{r}_{j}}\vec{F}_{j}
  + \vec{F}_{j}D^{\alpha}_{\vec{r}_{j}}\rho(x^{N};t).
 \]
 If $\vec{F}_{j}$ does not depend on $\vec{p}_{j}$,
 and $\vec{v}_{j}$ does not depend on $\vec{r}_{j}$,
 we get
 \[
  \frac{\partial }{\partial t}\rho(x^{N};t)
  +
  \sum^{N}_{j=1}\vec{v}_{j}D^{\alpha}_{\vec{r}_{j}}\rho(x^{N};t)+\sum^{N}_{j=1}\vec{F}_{j}D^{\alpha}_{\vec{p}_{j}}
  \rho(x^{N};t)=0,
 \]
 \[
  \vec{v}_{j}=D^{\alpha}_{\vec{p}_{j}}H(\vec{r},\vec{p}),  \vec{F}_{j}=-D^{\alpha}_{\vec{r}_{j}}H(\vec{r},\vec{p}) ,
 \]
 where $H(\vec{r},\vec{p})$ is a Hamiltonian of system with fractional derivatives.
 We get the Liouville equation in the form
 \begin{equation}\label{eq:2.3}
  \frac{\partial }{\partial t}\rho(x^{N};t)
  +
  \sum^{N}_{j=1}D^{\alpha}_{\vec{p}_{j}}H(\vec{r},\vec{p})D^{\alpha}_{\vec{r}_{j}}\rho(x^{N};t)
  -
  \sum^{N}_{j=1}D^{\alpha}_{\vec{r}_{j}}H(\vec{r},\vec{p})D^{\alpha}_{\vec{p}_{j}}\rho(x^{N};t)
  =0,
 \end{equation}
 or
 \begin{equation}\label{eq:2.4}
  \frac{\partial }{\partial t}\rho(x^{N};t)+iL_{\alpha}\rho(x^{N};t)=0,
 \end{equation}
 where $iL_{\alpha}$ is the Liouville operator with the fractional derivatives,
 \begin{equation}\label{eq:2.5}
  iL_{\alpha}\rho(x^{N};t)
  =
  \left(\sum^{N}_{j=1}D^{\alpha}_{\vec{p}_{j}}H(\vec{r},\vec{p})D^{\alpha}_{\vec{r}_{j}}
  -
  \sum^{N}_{j=1}D^{\alpha}_{\vec{r}_{j}}H(\vec{r},\vec{p})D^{\alpha}_{\vec{p}_{j}}\right)\rho(x^{N};t).
 \end{equation}

 A solution the Liouville equation (\ref{eq:2.5}) will be found
 with the Zubarev nonequilibrium statistical operator method~\cite{Zubarev19811509,Zubarev20021}.
 After choosing parameters of the reduced description,
 taking into account projections
 we present the non-equilibrium particle function $\rho\left(x^{N};t\right)$
 (as a solution of the Liouville equation) in general form
 \begin{equation}\label{eq:2.6}
   \rho(x^{N};t)
   =
   \rho_{rel}\left(x^{N};t\right)
   -
   \int^{t}_{-\infty}e^{\varepsilon(t'-t)}T(t,t')(1-P_{rel}(t'))iL_{\alpha}\rho_{rel}(x^{N};t')dt',
 \end{equation}
 where
 $T(t,t')=\exp_{+}\left(-\int^{t}_{t'}(1-P_{rel}(t'))iL_{\alpha}dt'\right)$
 is the evolution operator containing the projection,
 $\exp_+$ is ordered exponential,
 $\varepsilon\to+0$ after taking the thermodynamic limit,
 $P_{rel}(t')$ is the generalized Kawasaki-Gunton projection operator
 depended on a structure of the relevant statistical operator (distribution function), $\rho_{rel}(x^{N};t')$.
 By using the Zubarev nonequilibrium statistical operator method~\cite{Zubarev19811509,Zubarev20021,Zubarev20022} and approach~\cite{Markiv2011785},
 $\rho_{rel}\left(x^{N};t'\right)$ will be found from the extremum of the Renyi entropy
 at fixed values of observed values $\langle \hat{P}_{n}(x)\rangle^{t}_{\alpha}$,
 taking into account the normalization condition $\langle 1 \rangle^{t}_{\alpha, rel}=1$,
 where the nonequilibrium average values are found respectively~\cite{Tarasov2005082106,Tarasov2006052107,Tarasov2005011102,Tarasov2006033108}
 \begin{equation}\label{eq:2.7}
  \langle \hat{P}_{n} (x)\rangle^{t}_{\alpha}=\hat{I}^{\alpha}(1,\ldots,N)\hat{T}(1,\ldots,N)\hat{P}_{n}\rho(x^{N};t).
 \end{equation}
 $\hat{I}^{\alpha}(1,\ldots,N)$ has the following form for a system of $N$ particles
 \[
  \hat{I}^{\alpha}(1,\ldots,N)
  =
  \hat{I}^{\alpha}(1),\ldots,\hat{I}^{\alpha}(N),
  \quad
  \hat{I}^{\alpha}(j)=\hat{I}^{\alpha}(\vec{r}_{j})\hat{I}^{\alpha}(\vec{p}_{j})
 \]
 and defines operation of integration
 \begin{equation}\label{eq:2.8}
  \hat{I}^{\alpha}(x)f(x)=\int^{\infty}_{-\infty}f(x)d\mu_{\alpha}(x),
  \quad
  d\mu_{\alpha}(x)=\frac{|x|^{\alpha}}{\Gamma (\alpha)}dx.
 \end{equation}
 The operator $\hat{T}(1,\ldots,N)=\hat{T}(1),\ldots,\hat{T}(N)$ defines the operation
 \[
  \hat{T}(x_{j})f(x_{j})=\frac{1}{2} \left(f(\ldots,x'_{j}-x_{j},\ldots)+f(\ldots,x'_{j}+x_{j},\ldots)\right).
 \]
 Accordingly,
 the average value,
 calculated with the relevant distribution function,
 is defined as
 \[
 \langle (\ldots) \rangle^{t}_{\alpha, rel}=\hat{I}^{\alpha}(1,\ldots,N)\hat{T}(1,\ldots,N)(\ldots)\rho_{rel}(x^{N};t).
 \]
 According to~\cite{Markiv2011785},
 the relevant distribution function has the form
 \begin{equation}\label{eq:2.9}
  \rho_{rel}(t)
  =
  \frac{1}{Z_{R}(t)}
  \left(1
  -
  \frac{q-1}{q}
  \beta
  \left(
  H-\sum_{n}\int d\mu_{\alpha}(x)F_{n}(x;t)\delta \hat{P}_{n}(x;t)
  \right)
  \right)^{\frac{1}{q-1}},
 \end{equation}
 where $Z_{R}(t)$ is the partition function of the Renyi distribution,
 which is determined from the normalization condition and has the form
 \begin{align}\label{eq:2.10}
   Z_{R}(t)&=\hat{I}^{\alpha}(1,\ldots,N)\hat{T}(1,\ldots,N) \nonumber\\
           &\quad \times
            \left(
            1-\frac{q-1}{q}\beta\left(H-
 \sum_{n}\int d\mu_{\alpha}(x)F_{n}(x;t)\delta \hat{P}_{n}  (x;t)\right)
            \right)^{\frac{1}{q-1}}.
 \end{align}
 The parameters $F_{n}(x;t)$ are determined from the self-consistency conditions
 \begin{equation}\label{eq:2.11}
   \langle \hat{P}_{n}(x) \rangle^{t}_{\alpha}= \langle \hat{P}_{n}(x) \rangle^{t}_{\alpha, rel}.
 \end{equation}

 In the next section we will consider a specific example of diffusion processes in dense gases and liquids.

\section{Generalized diffusion equation with fractional derivatives}

 The nonequilibrium particle number density
 $n(\vec{ r}_{\alpha};t)=\langle\hat n(\vec{r})\rangle^{t}_{\alpha}$
 (where $n(\vec{r})=\sum_{j=1}^{N}\delta(\vec r-\vec r_{j})$ is the microscopic particle number density)
 is the main parameter of the reduced description
 for describing diffusion processes in classical gases and liquids.
 For such choice of parameter of the reduced description,
 the relevant distribution function has the form
 \begin{equation}\label{eq:2.12}
  \rho_{rel}(t)
  =
  \frac{1}{Z_{R}(t)}
  \left(
  1-\frac{q-1}{q}\beta
  \left(H-\int d\mu_{\alpha}(\vec r)\nu (\vec r;t)\delta\hat n (\vec r_{\alpha};t)\right)
  \right)^{\frac{1}{q-1}},
 \end{equation}
 where
 \begin{equation}\label{eq:2.13}
  Z_{R}(t)=\hat{I}^{\alpha}(1,\ldots,N)\hat{T}(1,\ldots,N)
  \left(1
  -\frac{q-1}{q}\beta
  \left(H-
  \int d\mu_{\alpha}(\vec r)\nu (\vec r;t)\delta\hat n (\vec r_{\alpha};t)\right)
  \right)^{\frac{1}{q-1}}
 \end{equation}
 is the partition function of the relevant distribution function,
 $\beta=\frac{1}{k_{\text{B}}T}$,
 $k_{\text{B}}$ is the Boltzmann constant,
 $T$ is the equilibrium value of temperature,
 $\delta\hat n(\vec r_{\alpha};t)=\hat n(\vec r)-\left\langle\hat n(\vec r)\right\rangle^{t}_{\alpha}$
 is the fluctuations of the density,
 and parameter $\nu(\vec r;t)$ is determined by the self-consistency condition
 \begin{equation}\label{eq:2.14}
  \left\langle \hat n(\vec{ r})\right\rangle^{t}_{\alpha}=\left\langle \hat n(\vec{ r})\right\rangle^{t}_{\alpha, rel}.
 \end{equation}
 It should be noted
 that the relevant distribution function (\ref{eq:2.12}) within the Renyi statistics for $q=1$
 coincides with one within the Gibbs statistics~\cite{Zubarev19811509}.
 The distribution (\ref{eq:2.12}) can be represented as
 \begin{equation}\label{eq:2.15}
   \rho_{rel}(t)
   =
   \frac{1}{Z_{R}(t)}
   \left(
    1-\frac{q-1}{q}\beta
    \left(
      H-\int d\mu_{\alpha}(\vec r)\nu^{*} (\vec r;t)\hat n (\vec r)
    \right)
   \right)^{\frac{1}{q-1}},
 \end{equation}
 where
 \begin{equation}\label{eq:2.16}
  Z_{R}(t)=\hat{I}^{\alpha}(1,\ldots,N)\hat{T}(1,\ldots,N)
  \left(
  1-\frac{q-1}{q}\beta
  \left(H-\int d\mu_{\alpha}(\vec r)\nu^{*} (\vec r;t)\hat n (\vec r)\right)
  \right)^{\frac{1}{q-1}},
 \end{equation}
 \[
   \nu^{*}(\vec r;t)
   =
   \frac{\displaystyle\nu(\vec r;t)}
   {\displaystyle 1+\frac{q-1}{q}\int d\mu_{\alpha}(\vec r) \nu(\vec r;t)\left\langle\hat n(\vec r)\right\rangle^{t}_{\alpha}}.
 \]
 Substituting Eq.~(\ref{eq:2.15}) in Eq.~(\ref{eq:2.6}),
 we find the nonequilibrium statistical operator
 \begin{equation}\label{eq:2.17}
  \rho(t)=\rho_{rel}(t)
  +
  \int_{-\infty}^{t}e^{\varepsilon(t'-t)}
  T(t,t')
  \int d\mu_{\alpha}(\vec r')I_{n}(\vec r'_{\alpha};t')\rho_{rel}(t)\nu^{*}(\vec r';t)dt',
 \end{equation}
 where
 \begin{equation}\label{eq:2.18}
  I_{n}(\vec r_{\alpha};t)=\left(1-P(t)\right)\frac{1}{q}\psi^{-1}(t)iL_{\alpha}\hat n(\vec r)
 \end{equation}
 is the generalized flow,
 \[
  \psi(t)=1-\frac{q-1}{q}\beta\left(H-\displaystyle\int d\mu_{\alpha}(\vec r)\nu^{*}(\vec    r;t)\hat n(\vec r)\right).
 \]
 $P(t)$ is the projection operator that has the following structure:
 \[
  P(t)\ldots=\int d\mu_{\alpha}(\vec r) \int d\mu_{\alpha}(\vec r')
  \left\langle\ldots\hat n (\vec r)\right\rangle_{\alpha, rel}^{t}
  [\langle\hat n(\vec r)\delta\{[q\psi(t)]^{-1}\hat n(\vec r')\}\rangle_{\alpha,rel}^{t}]^{-1}
  \delta\{[q\psi(t)]^{-1}\hat n (\vec r')\},
 \]
 where
 $\delta\{A\}=A-\langle A\rangle_{\alpha, rel}^{t}$.

 By using the nonequilibrium statistical operator (\ref{eq:2.17}),
 we get the generalized diffusion equation
 for the parameter of the reduced description
 \begin{equation}\label{eq:2.19}
  \frac{\partial}{\partial t}\left\langle\hat n(\vec r)\right\rangle^{t}_{\alpha}=
  \int d\mu_{\alpha}(\vec r')
  \int_{-\infty}^{t}e^{\varepsilon(t'-t)}\phi_{nn}(\vec r, \vec r';t,t')\beta\nu^{*}(\vec r';t')dt',
 \end{equation}
 where
 \begin{align*}
   \varphi_{nn}(\vec r,\vec r';t,t')&=\hat{I}^{\alpha}(1,..,N)\hat{T}(1,..,N)iL_{\alpha}\hat n(\vec r)T(t,t')I_{n}(\vec r'_{\alpha};t')\rho_{rel}(x^{N};t')\\
    & =\frac{\partial^{\alpha} }{\partial \vec r^{\alpha}}\cdot D_{q}(\vec r,\vec           r';t,t') \cdot\frac{\partial^{\alpha} }{\partial\vec r^{'\alpha}}
 \end{align*}
 is the generalized transport kernel (memory function),
 in which the averaging is performed with the distribution function~(\ref{eq:2.15}).

 As a result,
 we get the non-Markovian diffusion equation with fractional derivatives
 \begin{equation}\label{eq:2.20}
   \frac{\partial}{\partial t}\left\langle\hat n(\vec r)\right\rangle^{t}_{\alpha}=\frac{\partial^{\alpha} }{\partial\vec r^{\alpha}}\cdot\displaystyle\int d\mu_{\alpha}(\vec r')\displaystyle\int_{-\infty}^{t}e^{\varepsilon(t'-t)}D_{q}(\vec r,\vec r';t,t')\cdot\frac{\partial^{\alpha} }{\partial\vec r ^{'\alpha}}\beta\nu^{*}(\vec r';t')dt',
 \end{equation}
 where
 \begin{align}\label{eq:2.21}
  D_{q}(\vec r, \vec r';t,t')&=\langle\hat{\vec{v}}(\vec r)T(t,t')\hat{\vec{v}}(\vec r')\rangle^{t}_{\alpha, rel}\nonumber\\
  &=\hat{I}^{\alpha}(1,\ldots,N)\hat{T}(1,\ldots,N)\nonumber\\
  &\quad\times\hat{\vec{v}}(\vec r)T(t,t')\hat{\vec{v}}\vec (r')
    \frac {1}{Z_{R}(t)}
    \left(1-\frac{q-1}{q}\beta\left(H-\int d\mu_{\alpha}(\vec r)\nu^{*}(\vec r;t)\hat n(\vec r)\right)\right)^{\frac{1}{q-1}}
 \end{align}
 is the generalized diffusion coefficient within the Renyi statistics,
 in which the averaging is performed with the distribution function (\ref{eq:2.15}),
 where $\hat{\vec{v}}(\vec r)=\sum_{j=1}^{N}\vec v_{j}\delta(\vec r-\vec r_{j})$
 is the microscopic particle number flux density.
 At $q=1$
 the generalized diffusion equation within the Renyi statistics goes over to
 one within the Gibbs statistics with fractional derivatives.
 If $q=1$ and $\alpha=1$,
 we get the generalized diffusion equation within the Gibbs statistics~\cite{Zubarev19811509}.
 In the Markov approximation for the general diffusion coefficient in time and space
 $D_{q}(\vec r, \vec r';t,t')\approx D_{q}\delta (t-t')\delta (\vec r -\vec r')$,
 after exclusion the parameter $\nu^{*}(\vec r';t')$ by using the self-consistency condition,
 from Eq.~(\ref{eq:2.20}) we obtain the diffusion equation with fractional derivatives
 \begin{equation} \label{eq:2.22}
  \frac{\partial}{\partial t}\left\langle\hat n(\vec r)\right\rangle^{t}_{\alpha}
  =D_{q} \frac{\partial^{2\alpha} }{\partial r^{2\alpha}}
  \left\langle\hat n(\vec r)\right\rangle^{t}_{\alpha}.
 \end{equation}

 \section{Conclusion}

 By using the Zubarev nonequilibrium statistical operator~\cite{Zubarev19811509,Zubarev20021,Zubarev20022,Markiv2011785},
 the Liouville equation with fractional derivatives~\cite{Tarasov2005082106,Tarasov2006052107,Tarasov2005011102,Tarasov2006033108},
 and the principle of maximum entropy,
 we obtain the generalized (non-Markovian) diffusion equation with fractional derivatives.
 By using this approach,
 the generalized transfer equation with fractional derivatives can be obtained
 at some set of parameters of the reduced description $\langle\hat{P}_{n}(x)\rangle ^ {t}_{\ alpha}$
 of nonequilibrium state of the system
 In particular,
 if these parameters are nonequilibrium average values of particle density, momentum and energy,
 we obtain generalized hydrodynamic equations with fractional derivatives,
 which are generalization of Tarasov's results \cite{Tarasov2005286}.

\bibliography{MyBasa2}%
\bibliographystyle{apsrev} 

\end{document}